Consider the mixing of 200.0 mL of 0.200 M lanthanum nitrate, La(NO3 )3 , with 400.0 mL of 0.400 M potassium iodate, KIO3 . Calculate the concentration of each of the following ions in the solution: La3+, NO3 –, K+, and IO3 - .
Answer #1

La(NO3)3 + 3KIO3 I3LaO9 + 3 KNO3

Step1: Calculate the number of moles of each reactant ( before mixing)

Moles of La(NO3)3 = Molarity x volume ( litre)

= (0.200 x 200)/1000

= 0.04 moles

Moles of KIO3 = (0.400 x 400)/1000

= 0.16 moles

Step2: Calculate the concentration of reactant in the solution ( after mixing)

Total volume of solution = 200 + 400 = 600 ml

= 0.6 L

Molarity of La(NO3)3 = moles / volume

= 0.04/0.6

= 0.066 M

Molarity of KIO3 = 0.16/0.6

= 0.266 M

La(NO3)3 La+3 + 3NO3-

So, concentration of La+3 = 0.066 M

Concentration of NO3- = 3 x 0.066

= 0.198 M

KIO3 K+ + IO3-

So, concentration of KIO3 = concentration of K+ = concentration of IO3-

= 0.266 M